A modulating signal m(t)=10cos(2π×103t) is amplitude modulated with a carrier signal c(t)=50cos(2π×105t).

Find the modulation index, the carrier power, and the power required for transmitting AM wave.



1
Expert's answer
2021-12-12T16:41:01-0500

Solution;

"A_M=10"

"A_C=50"

Modulating index;

"M=\frac{A_M}{A_C}=\frac{10}{50}=0.2"

The carrier power;

"P_C=\frac12A_C^2=\frac{50^2}{2}=1250W"

Power required;

"P_R=P_C(1+\frac{M^2}{2})=1250(1+\frac{0.2^2}{2})=1275W"